# Pretraining LLM with Latent Thoughts in Continuous Space

### v1

@@ 4.1.1 Scaling Properties
As illustrated in Figure 1, our pretrained models demonstrate superior scaling properties in both parameter and data efficiency. Ours-1.26B (Pythia Arch), for instance, matches the performance of the official Pythia-2.8B with 55% fewer parameters. Furthermore, ours-1.4B (Pythia Arch) converges to the official version’s final performance using 62% less training data. Additional scaling curves on GPT-2 and LLaMA, presented in Appendix E, further validate the generalizability of our method.
@@ 4.1.2 Language Modeling Ability
To further quantify these pretraining gains, we evaluate perplexity (PPL) on several standard benchmarks (Pile validation, Wikitext (Merity et al., 2016), and the Lambada (Paperno et al., 2016)). The results in Figure 5 show that our method delivers substantial and consistent PPL reductions across all model sizes and datasets. Notably, ours-1.4B (Pythia Arch) is better than offical Pythia-2.8B.
@@ 4.2 Downstream Task Evaluation
Model (#training tokens)
Lambada
OpenAI
ARC
-E
Lambada
Standard
ARC
-C
Wino
Grande
PIQA
Hella
Swag
SciQ RACE
Avg acc /
@@
Ours-1.4B (Pythia Arch, 300B) 67.6 64.1 57.1 30.2 60.9 72.9 45.8 91.0 37.1 58.5 / +4.4
[…]
@@
Ours-1.4B (Pythia Arch, 300B) 63.6 67.4 56.0 32.6 64.0 73.5 46.4 94.5 37.9 59.5 / +5.4
We now evaluate the practical capabilities of our previous pretrained Pythia models on a range of downstream applications.
@@ 4.2.1 General Downstream Tasks
Datasets. Following (Gu & Dao, 2023; Zeng et al., 2025), we including LAMBADA (Paperno et al., 2016), SciQ (Welbl et al., 2017), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), ARC-Easy and ARC-Challenge (Clark et al., 2018), RACE (Lai et al., 2017) for comprehensive evaluation.
Baselines. We compare our pretrained Pythia models against several strong baselines: (1) the official Pythia models; (2) the PonderLM-Pythia models from prior work; (3) several open-source models trained on a similar data volume, including OPT (Zhang et al., 2022) and Bloom (Le Scao et al., 2023); and (4) TinyLLaMA (Zhang et al., 2024), a powerful model trained on ten times the amount of data (3T tokens vs. our 300B).
@@ 4.2.2 Instruction-Following Ability
We evaluate instruction-following capabilities by fine-tuning the 410m and 1.4B versions of ours (Pythia Arch) and the official Pythia models on the Alpaca dataset (Taori et al., 2023). When tested on the MT-Bench benchmark (Zheng et al., 2023), the our pretrained Pythia models consistently outperform their official Pythia counterparts across all categories, as shown in Figure 6. This results in significant average score improvements of 0.63 for the 410m model and 0.77 for the 1.4B model.
Figure 6: Instruction-following evaluation on MT-Bench. Our pretrained Pythia models outperform their official Pythia counterparts in all categories.
@@ 4.3 Comparison with Baseline Methods
Baselines. We compare our model against four strong approaches:
  • Looped Transformer (Saunshi et al., ): Processes the input by iterating through the entire set of transformer layers multiple times.
  • Pause Token (Goyal et al., 2023): Inserts a specified number of learnable “pause tokens” before generating each token, allowing for additional computation per step.
  • Pondering LLM (Zeng et al., 2025): Iteratively refines its output by feeding a “pondering embedding” (a probability-weighted sum of token embeddings) back into the model for several steps.
  • Scaled-up Model: As an oracle baseline, we train a standard LLaMA model with twice the number of parameters (2.8B). This model has inference FLOPs comparable to our method.
For the iterative baselines (Looped Transformer, Pause Token, and PonderLM), we evaluate them under two distinct computational budgets. First, we configure them to match the inference FLOPs of our method (a $2\times$ increase over the vanilla model), corresponding to 2 loops, 1 pause token, or 1 pondering step. Second, to provide a more challenging comparison, we pretrain them in a setting with double the inference FLOPs of our method (a $4\times$ increase), corresponding to 4 loops, 3 pause tokens, or 3 pondering steps.
Settings. We use the LLaMA-1.4B model as our testbed. For a fair comparison, all models are trained on a 26B token dataset using identical hyperparameters. We report perplexity (PPL; lower is better) on the Pile validation set, Wikitext, and the Lambada datasets (OpenAI and standard versions), as well as the average accuracy on the nine downstream tasks previously mentioned. The computational overhead is measured in relative inference FLOPs against the vanilla 1.4B model.
Results. The results, summarized in Section 4.3, show that our method consistently achieves the lowest perplexity across all language modeling benchmarks and the highest average accuracy on the downstream tasks. Notably, our approach not only surpasses all methods in the comparable ($2\times$) inference FLOPs category, including the LLaMA-2.8B oracle, but also demonstrates a significant advantage over methods operating at a much higher ($4\times$) inference budget. This highlights the superior performance and inference efficiency of our approach.
Table 3: Comparison on various benchmarks. Inference FLOPs are relative to the vanilla model. Actual throughput is evaluated following Wu & Tu (2024). Our method shows superior performance across all metrics while maintaining inference efficiency. Detailed downstream tasks performance are provided in Appendix D.
Model
Inference
FLOPs
Throughput
(tokens/s)
Pile
Lambada
OpenAI
Wikitext
Lambada
Standard
Avg Acc
0 shot
Avg Acc
5 shot
LLaMA-1.4B (train from scrach) $1\times$ 221.19 9.04 11.42 20.18 27.52 47.7 47.5
[…]
@@
LLaMA-2.8B (train from scrach) $2\times$ 110.50 8.23 8.93 18.09 17.08 49.8 50.6
Our LLaMA-1.4B $2\times$ 111.55 7.89 7.39 16.99 12.20 52.3 51.9
[…]
@@ 4.4 Effectiveness on off-the-shelf Foundation Models
We further investigate whether our method can effectively enhance existing, large-scale foundation models. To this end, we take the official LLaMA-3-3B (Grattafiori et al., 2024) model and perform continual pre-training on 5 billion tokens from the SlimPajama dataset (Shen et al., 2024). We compare the original model’s performance with two versions after this additional training: a vanilla continual pretraining baseline and our proposed approach. As illustrated in Figure 8, our method achieves a lower training loss than the vanilla baseline after consuming less than 1B tokens, and this performance gap widens as training progresses. As shown in Section 4.4, evaluation on our nine standard downstream tasks reveals that our method provides a substantial boost in performance, demonstrating its utility as a plug-and-play enhancement for off-the-shelf models.
Figure 7: Training loss comparison of our method and vanilla continual pretraining for the LLaMA-3-3B on 5B tokens from SlimPajama.
Figure 8: (Top) Impact of the num of Jacobi iterations . (Bottom) Chaining more latent thoughts before each token consistently lowers loss.
@@
Model
Lambada
OpenAI
ARC
-E
Lambada
Standard
ARC
-C
Wino
Grande
PIQA
Hella
Swag
SciQ
RACE
Avg acc /
@@
Ours CPT 69.0+2.2 78.6+0.5 67.4+3.3 48.0+3.9 72.9+1.5 78.1-0.5 57.0+0.9 96.7+0.3 41.4-0.4 67.7
@@ 4.5 Ablation Study
In this section, we study the impact of key components. We ablate the number of Jacobi iterations, different position embedding strategies, and the number of latent thoughts (chain analogous to CoT) generated before each token. All experiments are conducted on the Pythia-70m with 30B tokens.
As shown in Figure 8 (Top), increasing the number of Jacobi iterations initially lowers the loss, but the improvement saturates after only 3 iterations, which corroborates the fast convergence we observe (Figure 4). Meanwhile, as illustrated in Figure 8 (Bottom), chaining more latent thoughts consistently leads to better performance, which demonstrates our method’s potential for pretraining LLMs to generate a chain of latent thoughts before predicting each token. Regarding position embedding, we compared assigning sequential position ids to thoughts versus reusing the token’s position id for its corresponding thoughts and found a negligible performance difference. We use the latter strategy to avoid the need for long-context capabilities.
@@ 5 Conclusion
Appendix A Ethinic Statement
Our model’s demonstrated superiority in performance and efficiency necessitates a discussion of its dual-use nature. The same advanced capabilities that make it a powerful tool for positive applications could also be leveraged to generate highly convincing misinformation at scale. Furthermore, its novel architecture may present unknown security and privacy risks. We believe the responsible advancement of this technology requires a parallel effort to understand and mitigate these challenges.
Appendix B Reproducibility Statement
To ensure the reproducibility of our research, we have included the complete source code in the supplementary material. We also provide core hyperparameter and implementation settings in this paper. Using the provided code and specified configurations, all main results and figures reported in this paper can be fully reproduced. A public code repository will be made available upon publication.
Appendix C The Use of Large Language Models
Large Language Models (LLMs) were not used for the core methodology or the main research content of this paper. We utilized an LLM solely for the purpose of improving the language and clarity of the manuscript.
Appendix D Detailed Downstream Tasks Performance
@@
Model
Lambada
OpenAI
ARC
-E
Lambada
Standard
ARC
-C
Wino
Grande
PIQA
Hella
Swag
SciQ RACE
Avg acc /
@@
[…]
LLaMA-1.4B (train from scrach) 50.6 52.2 37.1 20.9 53.0 65.6 33.6 84.5 31.8 47.7
[…]
@@
LLaMA-2.8B (train from scrach) 54.3 53.4 43.5 24.4 53.1 68.0 36.2 83.4 31.5 49.8
Ours 58.1 58.0 48.2 25.2 53.9 70.7 38.6 85.9 32.4 52.3/+4.6
[…]
@@
[…]
LLaMA-1.4B (train from scrach) 45.1 53.5 34.7 22.3 50.9 66.1 33.6 89.3 31.6 47.5
[…]
@@
LLaMA-2.8B (train from scrach) 49.7 57.2 43.7 25.3 53.1 69.3 36.3 89.6 30.8 50.6
Ours 49.8 59.6 45.6 27.7 56.3 69.8 38.7 91.3 28.0 51.9 /+4.4
[…]
@@
Appendix E Generalization to GPT-2 and LLaMA
@@
Experimental Settings. We train both vanilla and latent-enhanced versions of these models from scratch on a subset of the Pile dataset, with sizes ranging from 405M to 1.4B parameters. The experimental setup, including the number of training tokens aligned with Chinchilla scaling laws (Hoffmann et al., 2022b). Detailed model configurations and training hyperparameters are provided in Table 6.
Results. The scaling curves are presented in Figure 9. Our method provides significant and consistent performance improvements for both GPT-2 and LLaMA across all model sizes. Notably, ours-834M (GPT-2 Arch) and ours-834M (LLaMA Arch) achieve a validation loss comparable to their vanilla counterparts trained with approximately 2.48x and 2.46x the parameter-token product, respectively.
@@
Parameters $\mathrm{n_{layers}}$ $\mathrm{d_{model}}$ $\mathrm{n_{heads}}$
Learning Rate
Batch Size
Training
Tokens

### v3

@@ 4.1.1 Scaling Properties
As illustrated in Figure 1, our pretrained models demonstrate superior scaling properties in both parameter and data efficiency. PonderLM-2-Pythia-1.26B, for instance, matches the performance of the official Pythia-2.8B with 55% fewer parameters. Furthermore, PonderLM-2-Pythia-1.4B converges to the official version’s final performance using 62% less training data. Additional scaling curves on GPT-2 and LLaMA, presented in Appendix B, further validate the generalizability of our method.
@@ 4.1.2 Language Modeling Ability
To further quantify these pretraining gains, we evaluate perplexity (PPL) on several standard benchmarks (Pile validation, Wikitext (merity2016pointer), and the Lambada (paperno2016lambada)). The results in Figure 4 show that our method delivers substantial and consistent PPL reductions across all model sizes and datasets. Notably, PonderLM-2-Pythia-1.4B is better than official Pythia-2.8B.
@@ 4.2 Downstream Task Evaluation
PonderLM-2-Pythia-1.4B (300B) 67.6 64.1 57.1 30.2 60.9 72.9 45.8 91.0 37.1 58.5 / +4.4
5-shot
@@
PonderLM-2-Pythia-1.4B (300B) 63.6 67.4 56.0 32.6 64.0 73.5 46.4 94.5 37.9 59.5 / +5.4
We now evaluate the practical capabilities of our previous pretrained Pythia models on a range of downstream tasks.
@@ 4.2.1 General Downstream Tasks
Datasets. Following (gu2023mamba; zeng2025pretraining), we including LAMBADA (paperno2016lambada), SciQ (welbl2017crowdsourcing), HellaSwag (zellers2019hellaswag), PIQA (bisk2020piqa), WinoGrande (sakaguchi2021winogrande), ARC-Easy and ARC-Challenge (clark2018think), RACE (lai2017race) for comprehensive evaluation.
[…]
@@ 4.2.2 Instruction-Following Ability
We assess instruction-following by fine-tuning the 410M and 1.4B versions of PonderLM-2-Pythia and the official Pythia models on Alpaca (alpaca). Evaluated on MT-Bench (zheng2023judging), our pretrained PonderLM-2-Pythia consistently outperforms the corresponding Pythia baselines across all categories (see Figure 5), improving the average score by 0.63 (410M) and 0.77 (1.4B).
Figure 5: Instruction-following evaluation on MT-Bench. Our PonderLM-2-Pythia models outperform their official Pythia counterparts in all categories.
@@ 4.3 Comparison with Baseline Methods
Table 3: Comparison on various benchmarks. Inference FLOPs are relative to the vanilla model. Actual throughput is evaluated following wu2024layercondensedkvcacheefficient. Our method shows superior performance across all metrics while maintaining inference efficiency. Detailed downstream tasks performance are provided in Appendix A.
Model Inference FLOPs Throughput (tokens/s) Pile Lambada OpenAI Wikitext Lambada Standard Avg Acc 0 shot Avg Acc 5 shot
LLaMA-1.4B (train from scratch) $1\times$ 221.19 9.04 11.42 20.18 27.52 47.7 47.5
Methods with comparable ($2\times$) inference FLOPs
@@
LLaMA-2.8B (train from scratch) $2\times$ 110.50 8.23 8.93 18.09 17.08 49.8 50.6
PonderLM-2-LLaMA-1.4B $2\times$ 111.55 7.89 7.39 16.99 12.20 52.3 51.9
Methods with higher inference FLOPs
@@
Pondering LLaMA-1.4B (7 steps) $8\times$ 26.39 8.01 7.94 17.21 14.23 51.4 51.5
Baselines. We compare against three compute-scaling baselines and a parameter-matched oracle: Looped Transformer (saunshireasoning), which repeats the full transformer stack for multiple loops; Pause Token (goyal2023think), which prepends a fixed number of learnable pause tokens before each generated token; Pondering LLM (zeng2025pretraining), which iteratively feeds back a probability-weighted “pondering embedding” for refinement; and a Scaled-up Model oracle, a standard LLaMA with doubled depth (doubling the number of layers to reach $\sim$2.8B parameters), whose inference FLOPs are comparable to our method. For iterative baselines (Looped Transformer, Pause Token, and PonderLM), we consider two compute budgets. First, we match our inference FLOPs (a $2\times$ increase over the vanilla model), using 2 loops, 1 pause token, or 1 pondering step. Second, for a stronger baseline, we pretrain them with substantially higher inference FLOPs than our method.
Settings. We use LLaMA-1.4B as the testbed. All models are trained on 26B tokens with identical hyperparameters for fair comparison. We report perplexity on Pile validation, WikiText, and LAMBADA (OpenAI and standard), along with average accuracy over the nine downstream tasks. Computational overhead is measured as relative inference FLOPs w.r.t. the vanilla 1.4B model.
Training Computation Analysis. We compare training FLOPs under the same data budget. Baselines with a $4\times$ inference budget (e.g., 4 loops / 3 pause tokens / 3 pondering steps) incur roughly $4\times$ training FLOPs, while the 2.8B scaled model costs about $2\times$ due to doubled parameters. Our training consists of: (i) one forward pass on the original sequence ($1\times$), (ii) $K$ Jacobi iterations on an interleaved token–thought sequence (doubling length, $2\times$ per iteration), and (iii) a final forward pass on the interleaved sequence ($2\times$). Thus the training FLOPs multiplier is $1+2K+2=3+2K$. With $K\in\{2,3\}$ and $\mathbb{E}[K]=2.5$, the average cost is $\approx 3+2\times 2.5=8\times$ the vanilla baseline.
Results. Training-compute lens. Although our method has higher training FLOPs when trained from scratch, it delivers better performance per training compute. Under comparable training FLOPs, we outperform PonderLM even with 7 pondering steps, and our 1.4B model achieves consistently higher mean downstream accuracy than TinyLLaMA-1.1B at matched training FLOPs despite TinyLLaMA using 33T tokens (Figure 1, right). Furthermore, this overhead can be largely mitigated by continual pretraining from existing checkpoints, where we observe consistent gains on off-the-shelf foundation models (Section 4.4). Inference-compute lens. The results in Table 3 further show that our method achieves the lowest perplexity across all language modeling benchmarks and the highest average downstream accuracy under matched inference FLOPs. Notably, at the comparable $2\times$ inference budget, we surpass all baselines including the $2\times$-parameter (doubled-depth) LLaMA-2.8B oracle, and we also maintain a clear advantage over methods operating at higher ($4\times$ and beyond) inference budgets.
@@ 4.4 Effectiveness on off-the-shelf Foundation Models
We further examine whether our method can improve existing large foundation models. Specifically, we start from the official LLaMA-3-3B (llama3herd2024) and conduct continual pre-training on 5B tokens from SlimPajama (shen2024slimpajamadc). We compare the original model with a vanilla continual pre-training baseline and our approach. As shown in Figure 7, our method attains lower training loss after consuming fewer than 1B tokens, and the gap continues to grow with training. On our nine standard downstream tasks (Table 4), it yields a clear performance gain, highlighting its plug-and-play value for off-the-shelf models.
Figure 6: Training loss of our method and vanilla continual pretraining for the LLaMA-3-3B on 5B tokens from SlimPajama.
Figure 7: (Top) Impact of the num of Jacobi iterations . (Bottom) Chaining more latent thoughts before each token lowers loss.
@@
PonderLM-2 CPT 69.0+2.2 78.6+0.5 67.4+3.3 48.0+3.9 72.9+1.5 78.1-0.5 57.0+0.9 96.7+0.3 41.4-0.4 67.7
4.5 Complementing with Test-Time Scaling
We verify whether our method complements test-time scaling (TTS). We reuse the two continual-pretrained models from Section 4.4: Standard CPT and PonderLM-2 CPT, both initialized from LLaMA-3-3B and trained on the same SlimPajama stream. We evaluate on GSM8K using Exact Match (EM). For TTS, we apply Majority Voting and Best-of-$N$ with $N\in\{1,\dots,10\}$ samples, and additionally test Chain-of-Thought (CoT) prompting. Figure 8 shows that both Voting and Best-of-$N$ improve as $N$ increases, while PonderLM-2 CPT consistently achieves higher EM and typically obtains larger gains than Standard CPT. Table 5 further shows that CoT boosts both models, with PonderLM-2 CPT benefiting more, indicating that our method is complementary to explicit reasoning prompting.
Figure 8: Exact Match (EM) under two TTS strategies: Majority Voting (left) and Best-of-$N$ (right), with $N\in\{1,\dots,10\}$ sampled solutions. EM increases with NN for both methods, while PonderLM-2 CPT consistently outperforms Standard CPT and typically yields larger gains.
4.6 Convergence Diagnostics of Jacobi Iteration
We evaluate the convergence of Jacobi updates in parallel training by tracking (i) the per-iteration change $r_{k}$ between successive iterates and (ii) the distance $d_{k}$ to the sequential autoregressive hidden states $H_{\text{seq}}$ (both measured by RMSE; see Figure 9). Convergence to a fixed point: As shown in Figure 9 (Left), $r_{k}$ drops rapidly and is approximately linear on a semi-log scale in early iterations, indicating near-exponential decay, before saturating at the BFloat16 numerical floor. Agreement with sequential inference: Figure 9 (Right) shows that $d_{k}$ reaches the same floor within a few iterations, suggesting that the Jacobi iterates match the sequential solution up to numerical precision.
Figure 9: Empirical convergence of Jacobi iteration on our pretrained model. (Left) RMSE between adjacent iterates $r_{k}=\mathrm{RMSE}(H^{(k)}-H^{(k-1)})$ rapidly decays and then saturates at the BFloat16 floor. (Right) RMSE to sequential inference $d_{k}=\mathrm{RMSE}(H^{(k)}-H_{\text{seq}})$ quickly drops to the same floor, showing agreement with the sequential result. We report an empirical decay rate $L$ (the slope of a log-linear fit over early, pre-saturation iterations) and the corresponding goodness-of-fit $R^{2}$.
@@ 4.7 Ablation Study
In this section, we study the impact of key components. We ablate the number of Jacobi iterations, different position embedding strategies, and the number of latent thoughts (chain analogous to CoT) generated before each token. All experiments are conducted on Pythia-70m with 30B tokens.
As shown in Figure 7 (Top), increasing the number of Jacobi iterations initially lowers the loss, but the improvement saturates after only 3 iterations, which corroborates the fast convergence we observe
Table 5: CoT gains on GSM8K. Exact Match (%) with and without CoT for Standard CPT and PonderLM-2 CPT. CoT improves both models, and the improvement is larger for PonderLM-2 CPT.
Setting Standard CPT PonderLM-2 CPT
Without CoT 10.01 12.59
With CoT 24.26+14.25 32.90+20.31
(Figure 9). Meanwhile, as illustrated in Figure 7 (Bottom), chaining more latent thoughts consistently leads to better performance, which demonstrates our method’s potential for pretraining LLMs to generate a chain of latent thoughts before predicting each token. Regarding position embedding, we compared assigning sequential position ids to thoughts versus reusing the token’s position id for its corresponding thoughts and found a negligible performance difference. We use the latter strategy to avoid shrinking the context window.
@@ 5 Conclusion
6 Acknowledgements
This work is sponsored by the National Natural Science Foundation of China (NSFC) grant (No. 62576211) and the National Key Research and Development Program of China (No. 2023ZD0121402).
Appendix A Detailed Downstream Tasks Performance
@@
0-shot
LLaMA-1.4B (train from scratch) 50.6 52.2 37.1 20.9 53.0 65.6 33.6 84.5 31.8 47.7
Methods with comparable ($2\times$) inference FLOPs
@@
LLaMA-2.8B (train from scratch) 54.3 53.4 43.5 24.4 53.1 68.0 36.2 83.4 31.5 49.8
PonderLM-2-LLaMA-1.4B 58.1 58.0 48.2 25.2 53.9 70.7 38.6 85.9 32.4 52.3/+4.6
Methods with higher ($4\times$) inference FLOPs
@@
5-shot
LLaMA-1.4B (train from scratch) 45.1 53.5 34.7 22.3 50.9 66.1 33.6 89.3 31.6 47.5
Methods with comparable ($2\times$) inference FLOPs
@@
LLaMA-2.8B (train from scratch) 49.7 57.2 43.7 25.3 53.1 69.3 36.3 89.6 30.8 50.6
PonderLM-2-LLaMA-1.4B 49.8 59.6 45.6 27.7 56.3 69.8 38.7 91.3 28.0 51.9 /+4.4
Methods with higher ($4\times$) inference FLOPs
@@
Appendix B Generalization to GPT-2 and LLaMA
Figure 10: Scaling curve comparison: vanilla models against our PonderLM-2-GPT and PonderLM-2-LLaMA models.
@@ Appendix B Generalization to GPT-2 and LLaMA
Experimental Settings. We train both vanilla and latent-enhanced versions of these models from scratch on a subset of the Pile dataset, with sizes ranging from 405M to 1.4B parameters. The experimental setup, including the number of training tokens aligned with Chinchilla scaling laws (hoffmann2022empirical). Detailed model configurations and training hyperparameters are provided in Table 7.
Results. The scaling curves are presented in Figure 10. Our method provides significant and consistent performance improvements for both GPT-2 and LLaMA across all model sizes. Notably, PonderLM-2-GPT-834M and PonderLM-2-LLaMA-834M achieve a validation loss comparable to their vanilla counterparts trained with approximately 2.48x and 2.46x the parameter-token product, respectively.
@@
Parameters $\mathrm{n_{layers}}$ $\mathrm{d_{model}}$ $\mathrm{n_{heads}}$ Learning Rate Batch Size (tokens) Training Tokens